# Can we determine the exact size of the nucleon?: A comprehensive study of different radii

### v1

@@ II The charge and magnetic radii
Figure 4 (left panel) shows a comparison between the results obtained for the mean-square charge radius of the neutron from different experimental techniques (squares) Krohn and Ringo (1973); Koester et al. (1995); Kopecky et al. (1997); Atac et al. (2021a); Heacock et al. (2021), as described earlier, as well as the corresponding results from various analyses and phenomenological studies (circles) Atac et al. (2021b); Filin et al. (2020, 2021); Albloushi et al. (2024); Belushkin et al. (2007); Goharipour et al. (2024). The PDG value Navas et al. (2024) is represented by a vertical band. As can be seen, the results are in better agreement compared to the proton case. In the right panel, we compare the results obtained from different theoretical techniques, including lattice QCD (circles) Alexandrou et al. (2019, 2020); Shintani et al. (2019); Djukanovic et al. (2024a); Tsuji et al. (2024), holographic QCD (stars) Mamo and Zahed (2023); Xu et al. (2021); Ahmady et al. (2021), and chiral methods (up triangles) Liu et al. (2014); Yang and Wang (2020). Note that Trinhammer Trinhammer (2021) obtained by assuming the neutron structure to be of intrinsic origin, determined by a mass Hamiltonian on the intrinsic configuration space . Overall, the experimental, phenomenological, and theoretical results span the intervals fm2, fm2, and fm2, respectively. Their corresponding combined values are fm2, fm2, and fm2. Finally, combining all experimental, theoretical, and phenomenological results, we obtain fm2.
@@ III The axial charge radius
From past to present, there have been significant efforts to determine , , and as well as the dipole mass using experimental techniques Bernard et al. (2002); Cai et al. (2023); Ahrens et al. (1987); Gran et al. (2006); Park et al. (2012); Abe et al. (2015); Petti et al. (2024); Lyubushkin et al. (2009); Aguilar-Arevalo et al. (2010a, b); Butkevich and Perevalov (2014), data analyses and phenomenological studies Adamuscin et al. (2008); Bodek et al. (2008); Meyer et al. (2016); Hill et al. (2018); Alvarez-Ruso et al. (2019); Ünal et al. (2019); Kaiser and Weise (2024); Irani et al. (2023); Hashamipour et al. (2019); Tomalak (2021); Megias et al. (2020); Pate et al. (2024), and theoretical calculations Mondal et al. (2020); Xu et al. (2021); Sasaki and Yamazaki (2008); Yamazaki et al. (2009); Bratt et al. (2010); Capitani et al. (2012); Horsley et al. (2014); Bali et al. (2015); Green et al. (2014); Hasan et al. (2018); Bhattacharya et al. (2016); Alexandrou et al. (2017b, 2021a, 2021b, 2024a); Park et al. (2022); Ishikawa et al. (2018); Shintani et al. (2019); Ishikawa et al. (2021); Tsuji et al. (2024); Gupta et al. (2017); Jang et al. (2024); Djukanovic et al. (2022); Bali et al. (2020); Capitani et al. (2019); Green et al. (2017); Jang et al. (2020b); Chang et al. (2018); Alexandrou et al. (2024b); Meyer et al. (2022); Yao et al. (2017); Lutz et al. (2020); Hermsen et al. (2024); Atayev and Mamedov (2022); Mamedov and Nasibova (2023); Schindler et al. (2007); Flores-Mendieta and Hofmann (2006); Liu et al. (2018, 2023); Dahiya and Randhawa (2014); Wang et al. (2021a); Eichmann and Fischer (2012); Choi et al. (2010); Ramalho et al. (2025); Chen et al. (2021, 2022); Braun et al. (2006); Wang et al. (2006); Aliev and Savci (2007); Anikin et al. (2016); Chen et al. (2024); Chen (2024). From an experimental perspective, our knowledge of the axial form factors (AFFs) is more limited compared to the electromagnetic FFs. The primary sources of information come from two classes of experiments: (anti)neutrino scattering off protons or nuclei and charged pion electroproduction. Additionally, muon capture by a proton and parity-violating electron scattering provide alternative methods to probe the AFF at very low and higher momentum transfers, respectively. For neutrino scattering experiments, notable measurements include: the measurement of the double differential cross section for charged-current quasielastic (CCQE) scattering on carbon and the corresponding AFF by the MiniBooNE Collaboration Aguilar-Arevalo et al. (2010a); Butkevich and Perevalov (2014) and the first high-statistics measurement of the reaction by the MINERvA Collaboration Cai et al. (2023). For charged pion electroproduction, numerous measurements have been conducted over the decades. A particularly important and recent example is the measurements performed by the CLAS Collaboration Park et al. (2012) through the extraction of the pion-nucleon multipoles near the production threshold for the channel at relatively high momentum transfer. For a comprehensive review and discussion of these topics, see Refs. Bernard et al. (2002); Hashamipour et al. (2019) and references therein.
@@
The theoretical calculations of the axial properties of the nucleon, including , , , and , have been predominantly performed within the lattice QCD framework Alexandrou et al. (2017b, 2021a, 2021b, 2024a); Park et al. (2022); Ishikawa et al. (2018); Shintani et al. (2019); Ishikawa et al. (2021); Tsuji et al. (2024); Gupta et al. (2017); Jang et al. (2024); Djukanovic et al. (2022); Bali et al. (2020); Capitani et al. (2019); Green et al. (2017); Jang et al. (2020b); Yamazaki et al. (2009); Bratt et al. (2010); Capitani et al. (2012); Horsley et al. (2014); Bali et al. (2015); Green et al. (2014); Bhattacharya et al. (2016); Hasan et al. (2018); Sasaki and Yamazaki (2008); Chang et al. (2018); Alexandrou et al. (2024b); Meyer et al. (2022). However, significant investigations have also been conducted using other theoretical techniques, including holographic QCD Mondal et al. (2020); Xu et al. (2021); Atayev and Mamedov (2022); Mamedov and Nasibova (2023), chiral methods such as PT Yao et al. (2017); Lutz et al. (2020); Schindler et al. (2007); Flores-Mendieta and Hofmann (2006), the perturbative chiral quark model Liu et al. (2018, 2023), and the chiral constituent quark model Dahiya and Randhawa (2014); Wang et al. (2021a). Additionally, studies have employed DSEs approach Eichmann and Fischer (2012), the relativistic constituent quark model Choi et al. (2010); Ramalho et al. (2025), the quark+diquark model Chen et al. (2021, 2022), and LCSR Braun et al. (2006); Wang et al. (2006); Aliev and Savci (2007); Anikin et al. (2016). Recently, Chen et al. Chen et al. (2024) demonstrated that the slope of in the forward limit, as given in Eq. (2), does not represent the three-dimensional mean-square axial radius in the Breit frame but instead corresponds to a contribution to the mean-square spin radius.
@@ IV The mechanical and mass radii
From past to present, there have been significant efforts to study the mechanical properties of the nucleon Lorcé et al. (2019); Pagels (1966); Lorcé (2015); Teryaev (2016); Lowdon et al. (2017); Cotogno et al. (2019); Varma and Schweitzer (2020); Ji (2021); Ji and Liu (2022); Freese and Miller (2021); Cao et al. (2023); Won et al. (2024a); Yao et al. (2024b); Li et al. (2024); Burkert et al. (2018); Kumerički (2019); Burkert et al. (2023b); Goharipour et al. (2025); Mamo and Zahed (2020); Kharzeev (2021); Duran et al. (2023); Guo et al. (2021, 2023, 2024); Han et al. (2022); Wang et al. (2020, 2021b, 2021c, 2022a, 2022b, 2023a, 2023b, 2024a); Hatta et al. (2025); Mamo and Zahed (2022); Belitsky and Radyushkin (2005); Polyakov (2003); Shanahan and Detmold (2019a, b); Pefkou et al. (2022); Hackett et al. (2024); Wang et al. (2024b); Mamo and Zahed (2021); de Téramond et al. (2021); Fujita et al. (2022); Fujii et al. (2024); Anikin (2019); Azizi and Özdem (2020); Dehghan et al. (2025); Jung et al. (2014a, b); Goeke et al. (2007); Kim et al. (2021); Kim and Kim (2021); Won et al. (2022, 2024b); Cebulla et al. (2007); Kim et al. (2012); Garcia Martin-Caro et al. (2023); Neubelt et al. (2020); Owa et al. (2022); Neubelt et al. (2020); Owa et al. (2022); Chakrabarti et al. (2020); Choudhary et al. (2022); More et al. (2022); Nair et al. (2024); Cao et al. (2024); Kou and Chen (2023), including GFFs, internal forces, pressure, surface tension, mechanical and mass radii, geometric shape, and more. These studies have utilized a wide range of theoretical approaches and phenomenological frameworks. For phenomenological studies, one can refer to analyses based on experimental data from deeply virtual Compton scattering (DVCS) Burkert et al. (2018); Kumerički (2019); Burkert et al. (2023b); Goharipour et al. (2025) and near-threshold photo-production of , , and particles Mamo and Zahed (2020); Kharzeev (2021); Duran et al. (2023); Guo et al. (2021, 2023, 2024); Han et al. (2022); Wang et al. (2020, 2021b, 2021c, 2022a, 2022b, 2023a, 2023b, 2024a); Hatta et al. (2025), as well as holographic fits to lattice data Mamo and Zahed (2022) and studies based on skewness-dependent GPDs Belitsky and Radyushkin (2005); Polyakov (2003). In contrast, theoretical studies have employed a variety of models and approaches, including lattice QCD Shanahan and Detmold (2019a, b); Pefkou et al. (2022); Hackett et al. (2024); Wang et al. (2024b), holographic QCD Mamo and Zahed (2021); de Téramond et al. (2021); Fujita et al. (2022); Fujii et al. (2024), LCSR Anikin (2019); Azizi and Özdem (2020); Dehghan et al. (2025), various frameworks of soliton models Jung et al. (2014a, b); Goeke et al. (2007); Kim et al. (2021); Kim and Kim (2021); Won et al. (2022, 2024b) (particularly the chiral quark-soliton model Goeke et al. (2007); Kim et al. (2021); Kim and Kim (2021); Won et al. (2022, 2024b)), the Skyrme model Cebulla et al. (2007); Kim et al. (2012); Garcia Martin-Caro et al. (2023), the bag model Neubelt et al. (2020); Owa et al. (2022), the light-front quark-diquark model Chakrabarti et al. (2020); Choudhary et al. (2022), the light-front Hamiltonian approach More et al. (2022); Nair et al. (2024), data-driven dispersive approaches Cao et al. (2024), and configurational entropy methods Kou and Chen (2023).
@@ V Summary and conclusion
The proton charge radius has been a focal point of intense study, especially in light of the proton radius puzzle, which revealed discrepancies between measurements from muonic hydrogen spectroscopy and traditional electron scattering or hydrogen spectroscopy. Our investigation shows that muonic hydrogen spectroscopy yields smaller values ( fm) with remarkably high precision, while electron scattering and hydrogen spectroscopy results span a wider range. Overall, the experimental results cover the interval fm. The combined experimental value of fm aligns well with the phenomenological studies which their combination yield fm. Theoretical calculations, particularly from lattice QCD ( fm), tend to predict smaller values ( fm), highlighting the need for further refinement in both experimental and theoretical approaches. Combining all all experimental, theoretical, and phenomenological results, we obtain fm.
@@
The mechanical and mass radii provide unique insights into the spatial distribution of pressure, shear forces, and mass within the nucleon. For the proton mass radius , theoretical calculations ( fm) suggest smaller values compared to phenomenological studies ( fm). Similarly, for the mechanical radius , the combined theoretical and phenomenological values ( fm and fm, respectively) are slightly smaller than the corresponding mass radius, reflecting differences in the spatial distribution of pressure and mass within the nucleon. Combining all phenomenological and theoretical results yields fm and fm. Table I summarizes the combined experimental, lattice QCD, theoretical, and phenomenological results obtained in the present study for each radius separately. Additionally, the last column of the table includes the combination of all experimental, theoretical, and phenomenological results, providing a comprehensive overview of the nucleon’s electromagnetic, weak, and mechanical properties.
@@
This work has provided a comprehensive overview of the different nucleon radii, including the charge and magnetic radii, the axial radius, as well as the mechanical and mass radii, emphasizing the progress made and the challenges that remain. We have addressed the common but misleading interpretation of the proton radius as a simple measure of its size, underscoring the nuanced and context-dependent nature of nucleon radii. While significant advances have been achieved in both experimental and theoretical approaches, discrepancies such as the proton radius puzzle and the differences between theoretical and phenomenological results highlight the complexity of nucleon structure. Future experiments, such as those planned at Jefferson Lab, Mainz, and CERN, along with continued advancements in lattice QCD and other theoretical frameworks, will be crucial in achieving a unified understanding of the nucleon’s internal structure. By addressing these challenges, we can deepen our understanding of QCD and its role in shaping the fundamental properties of matter.

### v3

@@ II The charge and magnetic radii
Figure 4 (left panel) shows a comparison between the results obtained for the mean-square charge radius of the neutron from different experimental techniques (squares) Krohn and Ringo (1973); Koester et al. (1995); Kopecky et al. (1997); Atac et al. (2021a); Heacock et al. (2021), as described earlier, as well as the corresponding results from various analyses and phenomenological studies (circles) Atac et al. (2021b); Filin et al. (2020, 2021); Albloushi et al. (2024); Belushkin et al. (2007); Goharipour et al. (2025a). The PDG value Navas et al. (2024) is represented by a vertical band. As can be seen, the results are in better agreement compared to the proton case. In the right panel, we compare the results obtained from different theoretical techniques, including lattice QCD (circles) Alexandrou et al. (2019, 2020); Shintani et al. (2019); Djukanovic et al. (2024a); Tsuji et al. (2024), holographic QCD (stars) Mamo and Zahed (2023); Xu et al. (2021); Ahmady et al. (2021), and chiral methods (up triangles) Liu et al. (2014); Yang and Wang (2020). Note that Trinhammer Trinhammer (2021) obtained by assuming the neutron structure to be of intrinsic origin, determined by a mass Hamiltonian on the intrinsic configuration space . Overall, the experimental, phenomenological, and theoretical results span the intervals fm2, fm2, and fm2, respectively. Their corresponding combined values are fm2, fm2, and fm2. Finally, combining all experimental, theoretical, and phenomenological results, we obtain fm2.
@@ III The axial charge radius
From past to present, there have been significant efforts to determine , , and as well as the dipole mass using experimental techniques Bernard et al. (2002); Cai et al. (2023); Ahrens et al. (1987); Gran et al. (2006); Park et al. (2012); Abe et al. (2015); Petti et al. (2024); Lyubushkin et al. (2009); Aguilar-Arevalo et al. (2010a, b); Butkevich and Perevalov (2014), data analyses and phenomenological studies Adamuscin et al. (2008); Bodek et al. (2008); Meyer et al. (2016); Hill et al. (2018); Alvarez-Ruso et al. (2019); Ünal et al. (2019); Kaiser and Weise (2024); Irani et al. (2023); Hashamipour et al. (2019); Tomalak (2021); Megias et al. (2020); Pate et al. (2024), and theoretical calculations Mondal et al. (2020); Xu et al. (2021); Sasaki and Yamazaki (2008); Yamazaki et al. (2009); Bratt et al. (2010); Capitani et al. (2012); Horsley et al. (2014); Bali et al. (2015); Green et al. (2014); Hasan et al. (2018); Bhattacharya et al. (2016); Alexandrou et al. (2017b, 2021a, 2021b, 2024); Park et al. (2022); Ishikawa et al. (2018); Shintani et al. (2019); Ishikawa et al. (2021); Tsuji et al. (2024); Gupta et al. (2017); Jang et al. (2024); Djukanovic et al. (2022); Bali et al. (2020); Capitani et al. (2019); Green et al. (2017); Jang et al. (2020b); Chang et al. (2018); Alexandrou et al. (2025); Meyer et al. (2022); Bhattacharya et al. (2025); Tomalak et al. (2023); Yao et al. (2017); Lutz et al. (2020); Hermsen et al. (2024); Atayev and Mamedov (2022); Mamedov and Nasibova (2023); Schindler et al. (2007); Flores-Mendieta and Hofmann (2006); Liu et al. (2018, 2023); Dahiya and Randhawa (2014); Wang et al. (2021a); Eichmann and Fischer (2012); Ramalho and Tsushima (2016); Choi et al. (2010); Ramalho et al. (2025); Chen et al. (2021, 2022); Braun et al. (2006); Wang et al. (2006); Aliev and Savci (2007); Anikin et al. (2016); Chen et al. (2024); Chen (2024); Hermsen et al. (2025). From an experimental perspective, our knowledge of the axial form factors (AFFs) is more limited compared to the electromagnetic FFs. The primary sources of information come from two classes of experiments: (anti)neutrino scattering off protons or nuclei and charged pion electroproduction. Additionally, muon capture by a proton and parity-violating electron scattering provide alternative methods to probe the AFF at very low and higher momentum transfers, respectively. For neutrino scattering experiments, notable measurements include: the measurement of the double differential cross section for charged-current quasielastic (CCQE) scattering on carbon and the corresponding AFF by the MiniBooNE Collaboration Aguilar-Arevalo et al. (2010a); Butkevich and Perevalov (2014) and the first high-statistics measurement of the reaction by the MINERvA Collaboration Cai et al. (2023). For charged pion electroproduction, numerous measurements have been conducted over the decades. A particularly important and recent example is the measurements performed by the CLAS Collaboration Park et al. (2012) through the extraction of the pion-nucleon multipoles near the production threshold for the channel at relatively high momentum transfer. For a comprehensive review and discussion of these topics, see Refs. Bernard et al. (2002); Hashamipour et al. (2019) and references therein.
@@
The theoretical calculations of the axial properties of the nucleon, including , , , and , have been predominantly performed within the lattice QCD framework Alexandrou et al. (2017b, 2021a, 2021b, 2024); Park et al. (2022); Ishikawa et al. (2018); Shintani et al. (2019); Ishikawa et al. (2021); Tsuji et al. (2024); Gupta et al. (2017); Jang et al. (2024); Djukanovic et al. (2022); Bali et al. (2020); Capitani et al. (2019); Green et al. (2017); Jang et al. (2020b); Yamazaki et al. (2009); Bratt et al. (2010); Capitani et al. (2012); Horsley et al. (2014); Bali et al. (2015); Green et al. (2014); Bhattacharya et al. (2016); Hasan et al. (2018); Sasaki and Yamazaki (2008); Chang et al. (2018); Alexandrou et al. (2025); Meyer et al. (2022); Bhattacharya et al. (2025); Tomalak et al. (2023) (see Ref. Gupta (2024) for an overview). However, significant investigations have also been conducted using other theoretical techniques, including holographic QCD Mondal et al. (2020); Xu et al. (2021); Atayev and Mamedov (2022); Mamedov and Nasibova (2023), chiral methods such as PT Yao et al. (2017); Lutz et al. (2020); Schindler et al. (2007); Flores-Mendieta and Hofmann (2006), the perturbative chiral quark model Liu et al. (2018, 2023), and the chiral constituent quark model Dahiya and Randhawa (2014); Wang et al. (2021a). Additionally, studies have employed DSEs approach Eichmann and Fischer (2012), covariant spectator quark model Ramalho and Tsushima (2016), relativistic constituent quark model Choi et al. (2010); Ramalho et al. (2025), quark+diquark model Chen et al. (2021, 2022), and LCSR Braun et al. (2006); Wang et al. (2006); Aliev and Savci (2007); Anikin et al. (2016). Recently, Chen et al. Chen et al. (2024) demonstrated that the slope of in the forward limit, as given in Eq. (2), does not represent the three-dimensional mean-square axial radius in the Breit frame but instead corresponds to a contribution to the mean-square spin radius.
@@ IV The mechanical and mass radii
From past to present, there have been significant efforts to study the mechanical properties of the nucleon Lorcé et al. (2019); Pagels (1966); Lorcé (2015); Teryaev (2016); Lowdon et al. (2017); Cotogno et al. (2019); Varma and Schweitzer (2020); Ji (2021); Ji and Liu (2022); Freese and Miller (2021); Cao et al. (2023); Won et al. (2024a); Yao et al. (2024b); Li et al. (2024); Burkert et al. (2018); Kumerički (2019); Burkert et al. (2023b); Goharipour et al. (2025b); Mamo and Zahed (2020); Kharzeev (2021); Duran et al. (2023); Guo et al. (2021, 2023, 2024); Han et al. (2022); Wang et al. (2020, 2021b, 2021c, 2022a, 2022b, 2023a, 2023b, 2024a); Hatta et al. (2025); Mamo and Zahed (2022); Belitsky and Radyushkin (2005); Polyakov (2003); Shanahan and Detmold (2019a, b); Pefkou et al. (2022); Hackett et al. (2024); Wang et al. (2024b); Mamo and Zahed (2021); de Téramond et al. (2021); Fujita et al. (2022); Fujii et al. (2024); Sugimoto and Tsukamoto (2025); Anikin (2019); Azizi and Özdem (2020); Dehghan et al. (2025); Jung et al. (2014a, b); Goeke et al. (2007); Kim et al. (2021); Kim and Kim (2021); Won et al. (2022, 2024b, 2023); Cebulla et al. (2007); Kim et al. (2012); Garcia Martin-Caro et al. (2023); Fujii et al. (2025); Neubelt et al. (2020); Owa et al. (2022); Neubelt et al. (2020); Owa et al. (2022); Chakrabarti et al. (2020); Choudhary et al. (2022); More et al. (2022); Nair et al. (2024); Cao et al. (2024); Kou and Chen (2023); Broniowski and Ruiz Arriola (2025); Sain et al. (2025), including GFFs, internal forces, pressure, surface tension, mechanical and mass radii, geometric shape, and more. These studies have utilized a wide range of theoretical approaches and phenomenological frameworks. For phenomenological studies, one can refer to analyses based on experimental data from deeply virtual Compton scattering (DVCS) Burkert et al. (2018); Kumerički (2019); Burkert et al. (2023b); Goharipour et al. (2025b) and near-threshold photo-production of , , and particles Mamo and Zahed (2020); Kharzeev (2021); Duran et al. (2023); Guo et al. (2021, 2023, 2024); Han et al. (2022); Wang et al. (2020, 2021b, 2021c, 2022a, 2022b, 2023a, 2023b, 2024a); Hatta et al. (2025), as well as holographic fits to lattice data Mamo and Zahed (2022) and studies based on skewness-dependent GPDs Belitsky and Radyushkin (2005); Polyakov (2003). In contrast, theoretical studies have employed a variety of models and approaches, including lattice QCD Shanahan and Detmold (2019a, b); Pefkou et al. (2022); Hackett et al. (2024); Wang et al. (2024b), holographic QCD Mamo and Zahed (2021); de Téramond et al. (2021); Fujita et al. (2022); Fujii et al. (2024); Sugimoto and Tsukamoto (2025), LCSR Anikin (2019); Azizi and Özdem (2020); Dehghan et al. (2025), various frameworks of soliton models Jung et al. (2014a, b); Goeke et al. (2007); Kim et al. (2021); Kim and Kim (2021); Won et al. (2022, 2024b, 2023) (particularly the chiral quark-soliton model Goeke et al. (2007); Kim et al. (2021); Kim and Kim (2021); Won et al. (2022, 2024b, 2023)), the Skyrme model Cebulla et al. (2007); Kim et al. (2012); Garcia Martin-Caro et al. (2023); Fujii et al. (2025), the bag model Neubelt et al. (2020); Owa et al. (2022), the light-front quark-diquark model Chakrabarti et al. (2020); Choudhary et al. (2022), the light-front Hamiltonian approach More et al. (2022); Nair et al. (2024), data-driven dispersive approaches Cao et al. (2024), configurational entropy methods Kou and Chen (2023), meson dominance model Broniowski and Ruiz Arriola (2025), and light-front spectator model Sain et al. (2025).
@@ V Summary and conclusion
The proton charge radius has been a focal point of intense study, especially in light of the proton radius puzzle, which revealed discrepancies between measurements from muonic hydrogen spectroscopy and traditional electron scattering or hydrogen spectroscopy. Our investigation shows that muonic hydrogen spectroscopy yields smaller values ( fm) with remarkably high precision, while electron scattering and hydrogen spectroscopy results span a wider range. Overall, the experimental results cover the interval fm. The combined experimental value of fm aligns well with the phenomenological studies whose combination yields fm. Theoretical calculations, particularly from lattice QCD ( fm), tend to predict smaller values ( fm), highlighting the need for further refinement in both experimental and theoretical approaches. Combining all experimental, theoretical, and phenomenological results, we obtain fm.
@@
The mechanical and mass radii provide unique insights into the spatial distribution of pressure, shear forces, and mass within the nucleon. For the proton mass radius , theoretical calculations ( fm) suggest smaller values compared to phenomenological studies ( fm). Such a situation is also true for the combined theoretical and phenomenological values of the proton mechanical radius ( fm and fm, respectively). Combining all phenomenological and theoretical results yields fm and fm. The slightly smaller value of compared to reflect differences in the spatial distribution of pressure and mass within the nucleon. Table I summarizes the combined experimental, lattice QCD, theoretical, and phenomenological results obtained in the present study for each radius separately. Note that, in this review, we classify lattice QCD calculations as theoretical results. While Table I provides separate combined values for lattice QCD predictions (for readers interested specifically in these results), these are included in the overall theoretical averages presented in the “Theory” column. Additionally, the last column of the table includes the combination of all experimental, theoretical, and phenomenological results, providing a comprehensive overview of the nucleon’s electromagnetic, weak, and mechanical properties. The combined values together with their uncertainties represent simple averages of independent results, obtained via standard procedure explained before, for illustrative purposes only. A rigorous statistical combination would require accounting for data overlaps and theoretical systematic differences, which is beyond the scope of this review. We also note that theoretical predictions differ substantially in their underlying assumptions and methodologies. While our simple averaging provides a compact overview, detailed comparisons should consider these fundamental differences on a case-by-case basis.
@@
6.
Table 1 presents a summary of proton radius values and their quoted uncertainties from various approaches. We emphasize that determining a true world average with reliable uncertainty requires careful treatment of all systematic effects and correlations among different results, which is a considerable undertaking and not addressed here. Nevertheless, we have presented systematic comparisons of all available results in our figures. The apparent sub-percent uncertainties in some entries in Table 1 reflect only the precision within individual studies, not the global knowledge of the proton radius.
This work has provided a comprehensive overview of the different nucleon radii, including the charge and magnetic radii, the axial radius, as well as the mechanical and mass radii, emphasizing the progress made and the challenges that remain. We have addressed the common but misleading interpretation of the proton radius as a simple measure of its size, underscoring the nuanced and context-dependent nature of nucleon radii. We have tried to highlight the distinction between “radii” (quantified via form factors) and ambiguous “size” terminology often used in literatures. Considering the fact that the latter cannot even be properly defined in a field theory approach, a more focused inquiry from both experiment and theory is needed to address whether a universal nucleon radius exists. While significant advances have been achieved in both experimental and theoretical approaches, discrepancies such as the proton radius puzzle and the differences between theoretical and phenomenological results highlight the complexity of nucleon structure. Future experiments, such as those planned at Jefferson Lab, Mainz, and CERN, along with continued advancements in lattice QCD and other theoretical frameworks, will be crucial in achieving a unified understanding of the nucleon’s internal structure. By addressing these challenges, we can deepen our understanding of QCD and its role in shaping the fundamental properties of matter.